*****please note that the instructor modified questions on the spreadsheet so please read the problem and spreadsheet carefully*****. I'm a bit confused so please let me know the steps to get the answers as well as the answers. Thank you for your help!
Hip-Hop Co. manufactures and markets several products. Management is considering the future of one Problem 1 product, electronic keyboards, that has not been as profitable as planned. Since this product is manufac- CVP analysis and chan tured and marketed independently of the other products, its total costs can be precisely measured. Next P2 P3 years plans call for a $350 selling price per unit. The fixed costs for the year are expected to be $42,000, up to a maximum capacity of 700 units. Forecasted variable costs are $210 per unit. 8-3B Required 1. Estimate the keyboards break-even point in terms of (a) sales units and (b) sales dollars. 2. Prepare a CVP chart for keyboards like that in Exhibit 18.14. Use 700 keyboards as the maximum Check (1) Break-even 300 units number of sales units on the horizontal axis of the graph, and $250,000 as the maximum dollar amount on the vertical axis. 3. Prepare a contribution margin income statement showing sales, variable costs, and fixed costs for keyboards at the break-even point.
media%2Fd90%2Fd9071d47-9e14-4ae9-adae-47
Answer #1

Answer

  • All working forms part of the answer
  • Part 1

A

Sale price per unit

$                                         350.00

B

Variable cost per unit

$                                         210.00

C = A - B

Contribution margin per unit

$                                         140.00 Answer

D

Fixed costs

$                                   42,000.00

E = D/C

Break Even in Units

300 = Answer

A

Contribution margin per unit

$                                         140.00

B

Sale price per unit

$                                         350.00

C = (A/B) x 100

Contribution margin ratio

40% = Answer

D

Fixed costs

$                                   42,000.00

E =D/C

Break Even point in Dollars

$                                 105,000.00 = Answer

  • Part 2

A

Target Net Income

$                                   70,000.00

B

Fixed costs

$                                   42,000.00

C = A+B

Total Contribution margin required

$                                 112,000.00

D

Contribution margin per unit

$                                         140.00

E = C/D

Units required to at target income

800 units = ANSWER

  • Analyse Part

>NO, it would not be possible for Hip Hop to earn $ 70,000 in Income next year.

>This is because, no. of units required computed above is 800 units. However, the question states that maximum capacity is 700 units. Hence under normal circumstances 800 units cannot be produced and hence cannot earn $ 70,000 Net Income.